OpenStudy (anonymous):

Find domain of each of the following expressions sqrt 6-x

OpenStudy (hoblos):

6-x >=0 x <= 6

OpenStudy (anonymous):

Thankss

OpenStudy (hoblos):

welcome XD

OpenStudy (anonymous):

One question why did you changee the signn isnt it only change when you divide by a negative

OpenStudy (hoblos):

yup you are right and here i divided by a negative but i forgot to show this step 6-x >=0 -x >= -6 x<=6